Evaluate 11C6.
Huh???
Please Help!

Answer by stanbon(75887)

Evaluate 11C6.
---
= 11!/(11-6)!*6! = (11*10*9*8*7*6)/(1*2*3*4*5*6) = 462
----
That is the number of different groups of people in
a room with 11 people in it.
Note: A group is unique is it has one different person in it.
=================
Cheers,
Stan H.
=================